The coordinates of the ends of a focal chord of the parabola y2= 4ax are (x1, y1) and (x2, y2). Show that x1x2 = a2 and y1y2 = −4a2.


1
Expert's answer
2021-08-19T11:04:41-0400

Say point A (x1, y1) = "(ap_1^2, 2ap_1)" , and

point B (x2, y3) = "(ap_2^2, 2ap_2)" coordinates of the ends of a focal chord of the

parabola "y^2 = 4ax"


Then O, A and B are collinear, where O is the vertex.


Therefore, Slope of OA = Slope of OB


Or, "\frac{2ap_1}{(ap_1^2-a)} = \frac{2ap_2}{(ap_2^2-a)}"


On solving,, we get


"p_1p_2^2 - p_1 = p_1^2p_2 - p_2"


"p_1 p_2(p_2-p_1) = -(p_2-p_1)"


Or

"p_1 p_2 = -1"


Now

"x_1 x_2 = ap_1^2 . ap_2^2"

Or

"x_1 x_2 = a^2 (1) = a^2"

Therefore,

"x_1 x_2 = a^2 (Proved)"


Now

"y_1 y_2 = 2ap_1 2a p_2 = 4a^2p_1p_2 = -4a^2"

Therefore,

"y_1 y_2 = -4a^2 (Proved)"